# MAAdvisor: Zero-Shot Index Advisor using Multi-Agent LLMs

### v1

@@ 7.4. Qualitative Evaluation
An example of MAAdvisor’s planning steps and index recommendation on TPC-H 20% budget can be found in Figure 7. We can observe that MAAdvisor efficiently recommends the indexes within five sub-steps, including selection and combination. Since the index selected is just within the storage budget, the revision agent is not called by the planning agent. It is also important to note that MAAdvisor adopts the same storage cost estimation as the other baselines to increase efficiency. Therefore, the actual storage cost of the output is smaller than the estimation in the planning phase. MAAdvisor can also opt to construct and calculate the actual index storage cost at each sub-step at a cost of a higher algorithm runtime.
@@
In order to further analyze how well MAAdvisor addresses the index regression problem, we pick out the queries in each benchmark having any methods’ recommended indexes under 20% budget leading to a more than one-second’s regression. As shown in Figure 6, it is observed that DB2Advis has the most regression cases, trading off the high algorithm efficiency. Most heuristic-based advisors suffer from regressions on the most complicated DSB benchmark, since their optimization relies solely on the optimizer’s estimations. It is important to note that, at the workload level, significant improvements on some queries with a cost of regression on some other queries is acceptable as long as the overall execution time is decreased. On the other hand, although MAAdvisor does not result in the highest query time improvement on all mentioned queries, the revision agent manages to avoid significant regressions and therefore leads to a higher workload-level performance.
8. Ablation Studies
@@
In this section, we first conduct an ablation study to evaluate the effectiveness of MAAdvisor’s workload representation technique and multi-agent pipeline. We will then analyze MAAdvisor’s generalisation ability over the LLM backbone and workload scale.
8.1. Effectiveness of Workload Representation
@@
To validate the effectiveness of MAAdvisor’s workload representation step, we perform an ablation study in which the database schema and raw workload queries are directly input to the LLM agents. Instead of selecting from column candidates, the selection agent is now prompted to select one possible column index from the SQL queries. The result for this raw workload setting is presented in Table 6. We can observe that the performance drops significantly from 61.33% to 29.52%, which is reasonable since the planning and selection agents will plan and select indexes solely based on their domain knowledge and even without any demonstrations. In addition, the input context for the agents will be much longer, and LLMs’ hallucination and misunderstanding will increase significantly.
@@
| MAAdvisor (GPT-3.5-turbo) | 60.01 |
| MAAdvisor (QWen3 32b) | 60.34 |
| MAAdvisor (QWen3 14b) | 40.86 |
| MAAdvisor (DeepCoder 14b) | 40.27 |
| MAAdvisor (TPC-H 20x) | 61.31 |
8.2. Effectiveness of Multi-agent pipeline
To explore the effectiveness of our multi-agent pipeline, we prompt the LLM to directly select from the column candidates and construct indexes in one step. Given the column candidates and storage budget, LLM is called only once to output all the recommended indexes. We present the one-call recommendation result in Table 6.
8.3. Effectiveness of Regression Indicator
We further test the effectiveness of the specially trained regression indicator embedded in the revision agent. We apply the MAAdvisor to the TPC-H benchmark with the regression indicator removed. The revision agent will then only rely on the provided expert experience to revise the current index recommendation. We can observe that MAAdvisor’s performance only suffers from a small drop. The main reason is that for smaller budgets, the index recommendation with a multi-agent pipeline is straightforward without too many revisions required, as shown in Figure 7. In addition, the revision agent alone can handle most of the severely regressive indexes by analysing the index utility estimation and learning from the expert experiences.
8.4. Different LLM Backbones
To evaluate MAAdvisor’s performance given different LLM backbones, we adopt one LLM API GPT-3.5-turbo (Brown et al., 2020), and three open-source LLMs, namely Qwen3-32B (Yang et al., 2025), Qwen3-14B, and DeepCoder-14B (Luo et al., 2025). DeepCoder-14B is the latest LLM that is specially finetuned on code-related data and achieves SOTA performance in code-related tasks. The results of adopting different LLM backbones are shown in Table 6. Since MAAdvisor relies on the reasoning ability of the LLM, we can observe that GPT-3.5-turbo and Qwen3-32B have comparable performance to GPT-4o-0806 and outperform Qwen3-14B. Smaller LLMs can also perform index recommendation following our MAAdvisor pipeline, but with a lower performance. It is important to note that the DeepCoder, specially fine-tuned on code-related data, underperforms the Qwen3 model with the same parameter size, showing that the coding ability of LLMs is not a must, given the rich information provided by the column candidates. In contrast, the reasoning ability of the LLMs is more important.
8.5. Different Workload Scales
In real-world applications, the workloads may always be updating, and such a situation is usually referred to as a dynamic setting. In order to simulate the change of scales, we construct another TPC-H workload of 380 queries named TPC-H 20x, where 20 random queries are generated for each query template. As shown in Table 6, MAAdvisor’s performance is robust against different workload scales. In addition, since the number of column candidates for the TPC-H 20x is almost the same as TPC-H 1x, the algorithm runtime for TPC-H 20x only increases from an average of 88.80 seconds to 92.91 seconds.
@@ 9. Conclusion and Discussion
In this work, we present MAAdvisor, a novel zero-shot index advisor that leverages the reasoning power of large language models through a hierarchical multi-agent pipeline. By representing the workload using column candidates and decomposing the index recommendation task into coordinated sub-steps handled by specialized agents, MAAdvisor achieves superior performance, strong generalisability, and high efficiency without the need for database-specific training. Extensive experiments across diverse benchmarks demonstrate that MAAdvisor not only outperforms traditional heuristic and learning-based approaches but also surpasses existing prompt-based LLM methods.
It is also important to acknowledge the current limitations and outline potential directions for future work. The efficiency of MAAdvisor may be further improved through the design of more precise and effective instruction sets. Additionally, fine-tuning the agent reasoning pipeline (Wang et al., 2024a) offers a promising avenue to reduce planning overhead and facilitate deployment on local LLMs. Furthermore, by capitalizing on the advanced reasoning and generalization capabilities of LLMs, the proposed multi-agent framework could be extended to address the index recommendation problem across heterogeneous DBMS environments and diverse SQL dialects.
@@ Acknowledgements.
- He et al. (2016) Ji He, Jianshu Chen, Xiaodong He, Jianfeng Gao, Lihong Li, Li Deng, and Mari Ostendorf. 2016. Deep Reinforcement Learning with a Natural Language Action Space. In Proceedings of the 54th Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers), Katrin Erk and Noah A. Smith (Eds.). Association for Computational Linguistics, Berlin, Germany, 1621–1630. https://doi.org/10.18653/v1/P16-1153
@@
- Perera et al. (2021) R. Malinga Perera, Bastian Oetomo, Benjamin I. P. Rubinstein, and Renata Borovica-Gajic. 2021. DBA bandits: Self-driving index tuning under ad-hoc, analytical workloads with safety guarantees. In 37th IEEE International Conference on Data Engineering, ICDE 2021, Chania, Greece, April 19-22, 2021. IEEE, 600–611. https://doi.org/10.1109/ICDE51399.2021.00058
@@
- Shen et al. (2023) Yongliang Shen, Kaitao Song, Xu Tan, Dongsheng Li, Weiming Lu, and Yueting Zhuang. 2023. HuggingGPT: Solving AI Tasks with ChatGPT and its Friends in Hugging Face. In Advances in Neural Information Processing Systems, A. Oh, T. Naumann, A. Globerson, K. Saenko, M. Hardt, and S. Levine (Eds.), Vol. 36. Curran Associates, Inc., 38154–38180. https://proceedings.neurips.cc/paper_files/paper/2023/file/77c33e6a367922d003ff102ffb92b658-Paper-Conference.pdf
@@
- Siddiqui et al. (2022) Tarique Siddiqui, Wentao Wu, Vivek Narasayya, and Surajit Chaudhuri. 2022. DISTILL: low-overhead data-driven techniques for filtering and costing indexes for scalable index tuning. Proc. VLDB Endow. 15, 10 (June 2022), 2019–2031. https://doi.org/10.14778/3547305.3547309

### v2

@@ 7.4. Qualitative Evaluation
An example of AMAZe’s planning steps and index recommendation on TPC-H 20% budget can be found in Figure 6. We observe that AMAZe efficiently recommends the indexes within five sub-steps, including selection and combination. Since the selected index is within the storage budget, the revision agent is not called by the planning agent. Note that AMAZe adopts the same storage cost estimation as the other baselines. Therefore, the actual storage cost of the output is smaller than the estimation in the planning phase. AMAZe can also opt to construct and calculate the actual index storage cost at each sub-step at a cost of a higher algorithm runtime.
@@
In order to further analyze how well AMAZe addresses the index regression problem, we pick out the queries in each benchmark having any methods’ recommended indexes under 20% budget leading to a more than one second’s regression. As shown in Figure 7, it is observed that DB2Advis has the most regression cases, trading off the high algorithm efficiency. Most heuristic-based advisors suffer from regressions on the most complicated DSB benchmark, since their optimization relies solely on the optimizer estimations. It is important to note that, at the workload level, significant improvements on some queries with a cost of regression on some other queries is acceptable as long as the overall execution time is decreased. On the other hand, although AMAZe does not result in the highest query time improvement on all mentioned queries, the revision agent manages to avoid significant regressions and therefore leads to a higher workload-level performance.
7.5. Ablation Studies
@@
In this section, we first conduct an ablation study to evaluate the effectiveness of AMAZe’s workload representation technique and multi-agent pipeline. We will then analyze AMAZe’s generalisation ability over the LLM backbone and workload scale.
@@
Effectiveness of Workload Representation. To validate the effectiveness of AMAZe’s workload representation step, we perform an ablation study in which the database schema and raw workload queries are directly input to the LLM agents. Instead of selecting from column candidates, the selection agent is now prompted to select one possible column index from the SQL queries. The result for this raw workload setting is presented in Table 6. We can observe that the performance drops significantly from 61.33% to 29.52%, which is reasonable since the planning and selection agents will plan and select indexes solely based on their domain knowledge and even without any demonstrations. In addition, the input context for the agents will be much longer, and LLMs’ hallucination and misunderstanding will increase significantly.
Effectiveness of Multi-agent pipeline. To explore the effectiveness of our multi-agent pipeline, we prompt the LLM to directly select from the column candidates and construct indexes in one step. Given the column candidates and storage budget, LLM is called only once to output all the recommended indexes. We present the one-call recommendation result in Table 6.
Effectiveness of Regression Indicator. We further test the effectiveness of the specially trained regression indicator embedded in the revision agent. We apply the AMAZe to the TPC-H benchmark with the regression indicator removed. The revision agent will then only rely on the provided expert experience to revise the current index recommendation. We can observe that AMAZe’s performance only suffers from a small drop. The main reason is that for smaller budgets, the index recommendation with a multi-agent pipeline is straightforward without too many revisions required, as shown in Figure 6. In addition, the revision agent alone can handle most of the severely regressive indexes by analysing the index utility estimation and learning from the expert experiences.
@@
| AMAZe (GPT-3.5-turbo) | 60.01 |
| AMAZe (QWen3 32b) | 60.34 |
| AMAZe (QWen3 14b) | 40.86 |
| AMAZe (DeepCoder 14b) | 40.27 |
| AMAZe (TPC-H 20x) | 61.31 |
Different LLM Backbones To evaluate AMAZe’s performance given different LLM backbones, we adopt an API GPT-3.5-turbo (Brown et al., 2020), and three open-source LLMs, namely Qwen3-32b (Yang et al., 2025), Qwen3-14b and DeepCoder-14b (Luo et al., 2025). DeepCoder 14b is the latest LLM that is specially finetuned on code-related data and achieves SOTA performance in code-related tasks. The results of adopting different LLM backbones are shown in Table 6. Since AMAZe relies on the reasoning ability of the LLM, we observe that GPT-3.5-turbo and Qwen3-32b have performance comparable to GPT-4o-0806 and outperform Qwen3-14b. Smaller LLMs can also perform index recommendation following AMAZe pipeline, but with a lower performance. It is note-worthy that the DeepCoder, specially fine-tuned on code-related data, underperforms the Qwen3 model with the same parameter size, showing that the coding ability of LLMs is not a must, given the rich information provided by the column candidates. In contrast, the reasoning ability of LLMs is more important.
Different Workload Scales In real-world applications, the workloads may always be updating, and such a situation is usually referred to as a dynamic setting. In order to simulate the change of scales, we construct another TPC-H workload of 380 queries named TPC-H 20x, where 20 random queries are generated for each query template. As shown in Table 6, AMAZe’s performance is robust against different workload scales. In addition, since the number of column candidates for the TPC-H 20x is almost the same as TPC-H 1x, the algorithm runtime for TPC-H 20x only increases from an average of 88.80 seconds to 92.91 seconds.
@@ 8. Conclusion and Discussion
In this work, we present AMAZe, a novel zero-shot index advisor that leverages the reasoning power of large language models through a hierarchical multi-agent pipeline. By representing the workload using column candidates and decomposing the index recommendation task into coordinated sub-steps handled by specialized agents, AMAZe achieves superior performance, strong generalisability, and high efficiency without the need for database-specific training. Extensive experiments across diverse benchmarks demonstrate that AMAZe not only outperforms traditional heuristic and learning-based approaches but also surpasses existing prompt-based LLM methods.
It is also important to acknowledge the current limitations and outline potential directions for future work. The efficiency of AMAZe may be further improved through the design of more precise and effective instruction sets. Additionally, fine-tuning the agent reasoning pipeline (Wang et al., 2024a) offers a promising avenue to reduce planning overhead and facilitate deployment on local LLMs. Furthermore, by capitalizing on the advanced reasoning and generalization capabilities of LLMs, the proposed multi-agent framework could be extended to address the index recommendation problem across heterogeneous DBMS environments and diverse SQL dialects.
@@ Acknowledgements.
- Brucato et al. (2024) Matteo Brucato, Tarique Siddiqui, Wentao Wu, Vivek Narasayya, and Surajit Chaudhuri. 2024. Wred: Workload Reduction for Scalable Index Tuning. Proc. ACM Manag. Data 2, 1, Article 50 (March 2024), 26 pages. https://doi.org/10.1145/3639305
@@
- Chang et al. (2024) Zhuo Chang, Xinyi Zhang, Yang Li, Xupeng Miao, Yanzhao Qin, and Bin Cui. 2024. MFIX: An Efficient and Reliable Index Advisor via Multi-Fidelity Bayesian Optimization. In 2024 IEEE 40th International Conference on Data Engineering (ICDE). 4343–4356. https://doi.org/10.1109/ICDE60146.2024.00331
@@
- Kossmann et al. (2020) Jan Kossmann, Stefan Halfpap, Marcel Jankrift, and Rainer Schlosser. 2020. Magic mirror in my hand, which is the best in the land? an experimental evaluation of index selection algorithms. Proc. VLDB Endow. 13, 12 (July 2020), 2382–2395. https://doi.org/10.14778/3407790.3407832
@@
- Perera et al. (2021a) R. Malinga Perera, Bastian Oetomo, Benjamin I. P. Rubinstein, and Renata Borovica-Gajic. 2021a. DBA bandits: Self-driving index tuning under ad-hoc, analytical workloads with safety guarantees. In 37th IEEE International Conference on Data Engineering, ICDE 2021, Chania, Greece, April 19-22, 2021. IEEE, 600–611. https://doi.org/10.1109/ICDE51399.2021.00058
- Perera et al. (2021b) R. Malinga Perera, Bastian Oetomo, Benjamin I. P. Rubinstein, and Renata Borovica-Gajic. 2021b. DBA bandits: Self-driving index tuning under ad-hoc, analytical workloads with safety guarantees. In 2021 IEEE 37th International Conference on Data Engineering (ICDE). 600–611. https://doi.org/10.1109/ICDE51399.2021.00058
@@
- Siddiqui et al. (2022a) Tarique Siddiqui, Saehan Jo, Wentao Wu, Chi Wang, Vivek Narasayya, and Surajit Chaudhuri. 2022a. ISUM: Efficiently Compressing Large and Complex Workloads for Scalable Index Tuning. In Proceedings of the 2022 International Conference on Management of Data (Philadelphia, PA, USA) (SIGMOD ’22). Association for Computing Machinery, New York, NY, USA, 660–673. https://doi.org/10.1145/3514221.3526152
- Siddiqui et al. (2022b) Tarique Siddiqui, Wentao Wu, Vivek Narasayya, and Surajit Chaudhuri. 2022b. DISTILL: low-overhead data-driven techniques for filtering and costing indexes for scalable index tuning. Proc. VLDB Endow. 15, 10 (June 2022), 2019–2031. https://doi.org/10.14778/3547305.3547309
@@
- Wang and Yoneki (2024) Taiyi Wang and Eiko Yoneki. 2024. IA2: Leveraging Instance-Aware Index Advisor with Reinforcement Learning for Diverse Workloads. In Proceedings of the 4th Workshop on Machine Learning and Systems (Athens, Greece) (EuroMLSys ’24). Association for Computing Machinery, New York, NY, USA, 10–17. https://doi.org/10.1145/3642970.3655839
@@
- Wu et al. (2022) Wentao Wu, Chi Wang, Tarique Siddiqui, Junxiong Wang, Vivek Narasayya, Surajit Chaudhuri, and Philip A. Bernstein. 2022. Budget-aware Index Tuning with Reinforcement Learning. In Proceedings of the 2022 International Conference on Management of Data (Philadelphia, PA, USA) (SIGMOD ’22). Association for Computing Machinery, New York, NY, USA, 1528–1541. https://doi.org/10.1145/3514221.3526128